A parallel plate capacitor is made up two plates, each having an area of 8.0 x 10 ^-4 m² and separated from each other by 5.0 mm. Half of the space between the plate is filled with glass and other with mica. Find the capacitance of this capacitor.



Given: A = 8.0 x 10^-4m²



d = 5.0 mm = 0.005 m

1
Expert's answer
2022-02-08T10:02:13-0500

Given:

"A=\rm 8.0 * 10^{-4}\: m^2"

"d = \rm 5.0\: mm = 0.005\: m"

"\kappa_1=5"

"\kappa_2=6"


The capacitance of two parralel capacitors

"C=C_1+C_2=\frac{\epsilon_0k_1A/2}{d}+\frac{\epsilon_0k_1A/2}{d}"

"C=\frac{\epsilon_0A}{2d}(\kappa_1+\kappa_2)"

"C=\frac{8.85*10^{-12}* 8.0 * 10^{-4}}{2*0.005}(5+6)=7.8*10^{-12}\:\rm F"